# Three flavor QCD phase transition with Möbius domain wall fermions

### v1

@@ Acknowledgments
We would like to thank our collaborators in the JLQCD collaboration for their helpful discussions. This research benefitted from the computational resources made available by several institutions and projects. These include the Supercomputer Fugaku, provided by the RIKEN Center for Computational Science through the HPCI projects hp220233 and hp210032, as well as Usability Research ra000001. The work also utilized the Wisteria/BDEC-01 Supercomputer System at Tokyo University/JCAHPC through HPCI project hp220108, the Ito supercomputer at Kyushu University through HPCI projects hp190124 and hp200050, and the Hokusai BigWaterfall at RIKEN. This work was supported in part by JSPS KAKENHI grant Nos. 20H01907 and 21H01085. Additionally, we acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 "Strong-interaction matter under extreme conditions" – project number 315477589 – TRR 211. We also acknowledge the Grid Lattice QCD framework [^1] https://github.com/paboyle/Grid and its extension tailored for A64FX processors [24], which played a crucial role in generating the QCD configurations for this research. We thank N. Meyer and T. Wettig for their valuable discussions regarding the use of Grid for A64FX. For the measurements, we utilized the Bridge++[31] and Hadrons code[32], both of which significantly contributed to the success of this work.

### v2

@@ Acknowledgments
We would like to thank our collaborators in the JLQCD collaboration for their helpful discussions. This research benefitted from the computational resources made available by several institutions and projects. These include the Supercomputer Fugaku, provided by the RIKEN Center for Computational Science through the HPCI projects hp220233 and hp210032, as well as Usability Research ra000001. The work also utilized the Wisteria/BDEC-01 Supercomputer System at Tokyo University/JCAHPC through HPCI project hp220108, the Ito supercomputer at Kyushu University through HPCI projects hp190124 and hp200050, and the Hokusai BigWaterfall at RIKEN. This work was supported in part by JSPS KAKENHI grant Nos. 20H01907 and 21H01085. Additionally, author Yu Zhang acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 "Strong-interaction matter under extreme conditions" – project number 315477589 – TRR 211. We also acknowledge the Grid Lattice QCD framework [^1] https://github.com/paboyle/Grid and its extension tailored for A64FX processors [24], which played a crucial role in generating the QCD configurations for this research. We thank N. Meyer and T. Wettig for their valuable discussions regarding the use of Grid for A64FX. For the measurement, we utilized the Hadrons code [32], which significantly contributed to the success of this work.